# The lattice structure of negative Sobolev and extrapolation spaces

### v1

@@ 2. Abstract results on ordered Banach spaces
Step 2: The span is a monotonically complete Banach lattice. As noted in Step 1, the cone in $X^{\prime}$ is normal. Hence, $X^{\prime}_{+}$ is also normal with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X^{\prime}_{+}}$. Therefore, $\operatorname{span}X^{\prime}_{+}$ is an ordered Banach space with a normal cone and a vector lattice at the same time. For such spaces it is easy to check that they have an equivalent norm that turns them into a Banach lattice. To show that $\operatorname{span}X^{\prime}_{+}$ is monotonically complete, consider an increasing net $(x^{\prime}_{j})$ in $X^{\prime}_{+}$ that is norm-bounded in $\operatorname{span}X^{\prime}_{+}$. We need to show that it has a supremum in $\operatorname{span}X^{\prime}_{+}$. The net $(x^{\prime}_{j})$ is in particular norm-bounded in $X^{\prime}$ and hence, as it is increasing, it converges weak to a vector $x^{\prime}\in X^{\prime}_{+}$. It is easy to check that $x^{\prime}$ is the supremum of $(x^{\prime}_{j})$ in $X^{\prime}$ and hence in $\operatorname{span}X^{\prime}_{+}$.
@@
Step 3: The span is a KB-space if $X$ is reflexive. If $X$ is reflexive, so is $X^{\prime}$. Moreover, as noted above the cone of $X^{\prime}_{+}$ is normal. Therefore, every increasing norm-bounded net in $X^{\prime}_{+}$ is norm-convergent in $X^{\prime}$. [1, Theorem 2.45]. Applying Lemma 2.2, we immediately obtain that every increasing norm-bounded net in $X^{\prime}_{+}$ is also norm-convergent with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X^{\prime}_{+}}$. In other words, $\operatorname{span}X^{\prime}_{+}$ is a KB-space. ∎
@@
From Step 3 in the proof of Theorem 2.1, we see that the assumption $X$ is reflexive in Theorem 2.1 can be replaced with “every increasing norm-bounded net in $X^{\prime}_{+}$ is norm-convergent”.
This assumption is, for instance, satisfied if $X_{+}$ has non-empty interior, since then there exists an equivalent norm on $X^{\prime}$ that is additive on the positive cone, which in turn gives that every increasing norm-bounded net in $X^{\prime}_{+}$ is a Cauchy net.
@@
Let $X$ be a Banach lattice and let $Z$ be a reflexive ordered Banach space with a normal cone. Assume that there exists an operator $J\in\mathcal{L}(X,Z)_{+}$ and a sequence $(R_{n})\subseteq\mathcal{L}(Z,X)_{+}$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to the $\operatorname{id}_{Z}$ with respect to the weak operator topology. Then $\operatorname{span}(Z_{+})=Z_{+}-Z_{+}$ is a KB-space with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}Z_{+}}$.
@@
The dual space $\tilde{Z}:=X^{\prime}$ is a Banach lattice and the cone in $\tilde{X}:=Z^{\prime}$ is generating since the cone in $Z$ is normal [1, Theorem 2.40].
Moreover, the sequence $(R_{n}^{\prime}J^{\prime})\subseteq\mathcal{L}(Z^{\prime})$ converges to $\operatorname{id}_{Z^{\prime}}$ with respect to the weak-operator topology. Since $Z^{\prime}$, is reflexive, this topology coincides with the weak operator topology. So we can apply Theorem 2.1 to the spaces $\tilde{X}$ and $\tilde{Z}$ and the operators $J^{\prime}$ and $R_{n}^{\prime}$ to conclude that $\operatorname{span}\tilde{X}^{\prime}_{+}=\operatorname{span}Z_{+}$ is a KB-space. ∎
Even though surprising, $\operatorname{span}Z_{+}$ need not be reflexive in Corollary 2.4; see Example 4.2. This also shows that $\operatorname{span}X^{\prime}_{+}$ need not be reflexive in Theorem 2.1 even if $X$ is so.
A Banach lattice $X$ is said to have order continuous norm if every increasing and order-bounded net (equivalently, sequence) in $X_{+}$ is norm-convergent. Classical examples are $L^{p}$-spaces for $p\in[1,\infty)$ as well as the space $c_{0}$. KB-spaces have order continuous norm, but the converse is not true as the example $c_{0}$ shows. We refer to [14] for extensive literature on Banach lattices with order continuous norms.
@@
Let $X$ be a Banach lattice with order continuous norm and let $Z$ be an ordered Banach space. Assume that there exists a positive bijection $S\in\mathcal{L}(Z,X)$, a positive operator $J\in\mathcal{L}(X,Z)$ and a sequence of positive operators $(R_{n})\subseteq\mathcal{L}(Z,X)$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to the $\operatorname{id}_{Z}$ with respect to the strong operator topology.
Then $\operatorname{span}(Z_{+})$ is a Banach lattice with order continuous norm with respect to a norm that is equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}Z_{+}}$. If $X$ is a KB-space, then so is $\operatorname{span}(Z_{+})$.
@@
We point out that Theorem 2.5 gives that $\operatorname{span}Z_{+}$ is a Banach lattice if $X$ is a Banach lattice with order continuous norm, while Corollary 2.4 requires $X$ to be reflexive to get the same conclusion. Similarly, Theorem 2.5 gives that $\operatorname{span}Z_{+}$ is a KB-space Banach lattice if $X$ is so, while Corollary 2.4 again requires $X$ to be reflexive to get the same conclusion. The prize that one pays for this weaker assumption in Theorem 2.5 is that one also needs the existence of the bijective positive operator $S$ and that the sequence $(JR_{n})$ is required to converge with respect to the strong instead of the weak operator topology.
@@
Indeed, after replacing sequences with subsequences, we may assume that one has $\sum_{n=1}^{\infty}\left\lVert g_{n}-g\right\rVert_{Z}<\infty$. Then the values $\left\lVert Sg_{n}-Sg\right\rVert_{X}$ are also summable, so the series $h:=\sum_{n=1}^{\infty}\left\lvert Sg_{n}-Sg\right\rvert$ converges absolutely in $X$. One obtains that $0\leq Sf_{n}\leq Sg_{n}\leq h+Sg$ for all indices $n$, where the first two inequalities follow from the positivity of $S$ and the last inequality from the definition of $h$. So the sequence $(Sf_{n})$ is contained in an order interval of $X$ and since $X$ has order continuous norm, all order intervals in $X$ are weakly compact [14, Theorem 1.12]. Thus $(Sf_{n})$ has a subsequence that converges weakly in $X$. Since the inverse operator $S^{-1}:X\to Z$ (which might not be positive) is continuous by the bounded inverse theorem and is thus weakly continuous from $X$ to $Z$, it follows that $(f_{n})$ does indeed have a subsequence that converges weakly in $Z$.
Step 2: The span is a lattice. To see this, let $z\in\operatorname{span}Z_{+}$. It suffices to show that ±$\pm z$ has a supremum in $\operatorname{span}Z_{+}$. As $z\in Z_{+}-Z_{+}$, there exists $y\in Z_{+}$ such that ±$\pm z\leq y$. Then we have ±$\pm R_{n}z\leq R_{n}y$ and hence $0\leq J\left\lvert R_{n}z\right\rvert\leq JR_{n}y$ for all indices $n$. By assumption, $(JR_{n}y)$ converges to $y$ with respect to the norm in $Z$, and thus the first part of the proof shows that $(J\left\lvert R_{n}z\right\rvert)$ has a subsequence $(J\left\lvert R_{n_{j}}z\right\rvert)$ that converges weakly to an element $s\in Z$. As $Z_{+}$ is closed and convex it is weakly closed, so $s\in Z_{+}$. We show that $s$ is the supremum of ±$\pm z$ within $\operatorname{span}(Z_{+})$. For every upper bound $u\in\operatorname{span}(Z_{+})$ of ±$\pm z$ – repeating the above argument with $u$ instead of $y$ – we get
@@
Step 3: The span is a Banach lattice with order continuous norm. We first show that the cone $Z_{+}$ in $Z$ is normal. To see this, let $y,z\in Z$ satisfy $0\leq y\leq z$. Then we have $0\leq Sy\leq Sz$ in $X$ and hence, as $X$ is a Banach lattice, $\left\lVert Sy\right\rVert\leq\left\lVert Sz\right\rVert$. The inverse operator $S^{-1}$ is bounded by the bounded inverse theorem and thus,
@@
Therefore, $Z_{+}$ is normal in $Z$, and in turn, also in $\operatorname{span}Z_{+}$; see the discussion after (1.1). So $\operatorname{span}(Z_{+})$ is an ordered Banach space with a normal cone and, at the same time, a vector lattice. Hence, it is a Banach lattice with respect to an equivalent norm.
In order to show that the norm in $\operatorname{span}Z_{+}$ is order continuous, let $(z_{j})$ be an increasing order-bounded net in $Z_{+}$; we need to show that it is norm-convergent in $\operatorname{span}Z_{+}$. The net $(Sz_{j})$ in $X_{+}$ is also increasing and order-bounded, so it is norm-convergent in $X$, since $X$ has order continuous norm. By the bounded inverse theorem, we obtain that the net $(z_{j})$ is norm-convergent with respect to the norm in $Z$. As $(z_{j})$ was increasing, we may apply Lemma 2.2 to assert the claim.
Step 4: The span is a KB-space if $X$ is so. Since every increasing norm-bounded net in $\operatorname{span}(Z_{+})$ is also increasing and norm-bounded in $Z$, this is yet another application of Lemma 2.2. ∎

### v3

@@ 2. Abstract results on ordered Banach spaces
Step 2: The span is a monotonically complete Banach lattice. As noted in Step 1, the cone in $X^{\prime}$ is normal. Hence, $X^{\prime}_{+}$ is also normal with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(X^{\prime}_{+})}$. Therefore, $\operatorname{span}(X^{\prime}_{+})$ is an ordered Banach space with a normal cone and a vector lattice at the same time. Hence, Lemma 2.4 shows that there exists an equivalent norm that turns $\operatorname{span}(X^{\prime}_{+})$ into a Banach lattice. To show that $\operatorname{span}(X^{\prime}_{+})$ is monotonically complete, consider an increasing net $(x^{\prime}_{j})$ in $X^{\prime}_{+}$ that is norm-bounded in $\operatorname{span}(X^{\prime}_{+})$. We need to show that it has a supremum in $\operatorname{span}(X^{\prime}_{+})$. The net $(x^{\prime}_{j})$ is in particular norm-bounded in $X^{\prime}$ and hence, by Banach-Alaoglu, it has a subnet that converges weak to $x^{\prime}\in X^{\prime}_{+}$. Actually, since $(x^{\prime}_{j})$ is increasing, it even weak-converges itself to $x^{\prime}\in X^{\prime}_{+}$. It is easy to check that $x^{\prime}$ is the supremum of $(x^{\prime}_{j})$ in $X^{\prime}$ and hence in $\operatorname{span}(X^{\prime}_{+})$.
@@
(b) If $X$ is reflexive, so is $X^{\prime}$. Moreover, as noted above the cone of $X^{\prime}_{+}$ is normal. Therefore, every increasing norm bounded net in $X^{\prime}_{+}$ is norm convergent in $X^{\prime}$ [1, Theorem 2.45]. Applying Lemma 2.2, we immediately obtain that every increasing norm bounded net in $X^{\prime}_{+}$ is also norm convergent with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(X^{\prime}_{+})}$. In other words, $\operatorname{span}(X^{\prime}_{+})$ is a KB-space.
(c) We begin with the preliminary observations that $J$ is injective and bipositive and that $J^{\prime}$ is bipositive as well. Indeed, first note that the bipositivity of $J$ holds since if $x\in X$ and $Jx\geq 0$, then $x$ is the weak limit of the sequence $(R_{n}Jx)\subseteq X_{+}$ and hence, $x\geq 0$. Being a bipositive bounded operator between ordered Banach spaces, $J$ is injective [3, Proposition 2.1(a)]. To see the bipositivity of $J^{\prime}$, note that due to the assumption of (c), the net $(R_{n}^{\prime}J^{\prime})\subseteq\mathcal{L}(Z^{\prime})$ converges to $\operatorname{id}_{Z^{\prime}}$ with respect to the weak-operator topology. So if $z^{\prime}\in Z^{\prime}$ and $J^{\prime}z^{\prime}\geq 0$, then $z^{\prime}$ is positive since it is the weak-limit of $(R_{n}^{\prime}J^{\prime}z)$ and since $Z^{\prime}_{+}$ is weak-closed.
Next, we show that $J^{\prime}(Z^{\prime}_{+})$ is a face in $X^{\prime}_{+}$; as by Lemma 2.3 and the bipositivity of $J^{\prime}$ this implies (c). We roughly follow the arguments recently given in [3, Theorem 2.5] where a related result was shown. Let $z^{\prime}\in Z^{\prime}_{+}$ and $x^{\prime}\in X^{\prime}_{+}$ satisfy $0\leq x^{\prime}\leq J^{\prime}z^{\prime}$; we have to find a functional $y^{\prime}\in Z^{\prime}_{+}$ that satisfies $J^{\prime}y^{\prime}=x^{\prime}$ or, in other words, $\left\langle y^{\prime}\,,\,Jx\right\rangle=\left\langle x^{\prime}\,,\,x\right\rangle$ for all $x\in X$.
To this end consider the linear map $\varphi:JX\to\mathbb{R}$ that is given by $\varphi(z)=\left\langle x^{\prime}\,,\,J^{-1}z\right\rangle$ for all $z\in JX$. Observe that $\varphi$ is well-defined since $J$ is injective and it is positive since $J$ is bipositive and $x^{\prime}$ is positive. We intend to extend $\varphi$ to a functional in $y^{\prime}\in Z^{\prime}_{+}$; for such an extension $y^{\prime}$ to exist it is, according to the extension result in [3, Proposition 2.4], sufficient for $\varphi(v)\leq\langle z^{\prime},w\rangle$ to hold for all $v\in JX$ and all $w\in Z_{+}$ that satisfy $v\leq w$. So consider such $v$ and $w$ and observe that indeed
$\varphi(v)$ $=\left\langle x^{\prime}\,,\,J^{-1}v\right\rangle=\lim_{n}\left\langle x^{\prime}\,,\,R_{n}v\right\rangle$
$\leq\limsup_{n}\left\langle x^{\prime}\,,\,R_{n}w\right\rangle\leq\limsup_{n}\left\langle J^{\prime}z^{\prime}\,,\,R_{n}w\right\rangle=\left\langle z^{\prime}\,,\,w\right\rangle;$
the second equality uses the weak operator convergence of $(R_{n}J)$ to $\operatorname{id}_{X}$, the first inequality uses the positivity of $x^{\prime}$, the second inequality the positivity of $w$, and the last equality the weak operator convergence of $(JR_{n})$ to $\operatorname{id}_{Z}$. So [3, Proposition 2.4] gives us a functional $y^{\prime}\in Z^{\prime}_{+}$ that extends $\varphi$. The definition of $\varphi$ now readily gives that $J^{\prime}y^{\prime}=x^{\prime}$, so $J^{\prime}Z^{\prime}_{+}$ is indeed a face in $X_{+}$. ∎
@@
From the proof of Theorem 2.1(b), we see that the assumption that $X$ is reflexive in Theorem 2.1(b) can be replaced with the weaker assumption that every increasing norm bounded net in $X^{\prime}_{+}$ is norm convergent.
This assumption is, for instance, satisfied if $X_{+}$ has non-empty interior, since then there exists an equivalent norm on $X^{\prime}$ that is additive on the positive cone, which in turn gives that every increasing norm bounded net in $X^{\prime}_{+}$ is a Cauchy net.
@@
Let $X$ be a Banach lattice and let $Z$ be a reflexive ordered Banach space with a normal cone. Assume that there exists an operator $J\in\mathcal{L}(X,Z)_{+}$ and a sequence $(R_{n})\subseteq\mathcal{L}(Z,X)_{+}$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to $\operatorname{id}_{Z}$ with respect to the weak operator topology. Then $\operatorname{span}(Z_{+})=Z_{+}-Z_{+}$ is a KB-space with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(Z_{+})}$.
@@
The dual space $\tilde{Z}:=X^{\prime}$ is a Banach lattice. Moreover, for each $z\in Z$ we have $J(R_{n}z)^{+}-J(R_{n}z)^{-}\to z$ weakly, so the span of the cone $Z_{+}$ is dense in $Z$. Hence, the dual wedge $\tilde{X}_{+}=Z^{\prime}_{+}$ is actually a cone in $\tilde{X}:=Z^{\prime}$ and therefore $\tilde{X}$ is an ordered Banach space. Note that $\tilde{X}_{+}$ is generating because the cone in $Z$ is normal [1, Corollary 2.43].
Further, the sequence $(R_{n}^{\prime}J^{\prime})\subseteq\mathcal{L}(Z^{\prime})$ converges to $\operatorname{id}_{Z^{\prime}}$ with respect to the weak-operator topology. Since $Z^{\prime}$ is reflexive, this topology coincides with the weak operator topology. Applying Theorem 2.1(b) to the spaces $\tilde{X}$ and $\tilde{Z}$ and the operators $J^{\prime}$ and $R_{n}^{\prime}$, we conclude $\operatorname{span}(\tilde{X}^{\prime}_{+})=\operatorname{span}(Z_{+})$ is a KB-space. ∎
Even though surprising, $\operatorname{span}(Z_{+})$ need not be reflexive in Corollary 2.6; see Example 4.2. This also shows that $\operatorname{span}(X^{\prime}_{+})$ need not be reflexive in Theorem 2.1 even if $X$ is so.
Remark 2.7.
In the situation of Corollary 2.6 assume in addition that $X$ is reflexive and that $R_{n}J\to\operatorname{id}_{X}$ with respect to the weak operator topology. Then the sequence $(J^{\prime}R_{n}^{\prime})\subseteq\mathcal{L}(X^{\prime})$ also converges to $\operatorname{id}_{X^{\prime}}$ with respect to the weak operator topology. So it follows from Theorem 2.1(c) that $J$ is a lattice homomorphism from $X^{\prime\prime}=X$ to $\operatorname{span}(Z_{+})$ and that $JX$ is a lattice ideal in $\operatorname{span}(Z_{+})$.
A Banach lattice $X$ is said to have order continuous norm, if every increasing and order bounded net (equivalently, sequence) in $X_{+}$, is norm convergent. Classical examples are $L^{p}$-spaces for $p\in[1,\infty)$ as well as the space $c_{0}$. KB-spaces have order continuous norm, but the converse is not true as the example $c_{0}$ shows. We refer to [19] for a thorough and extensive presentation of the theory of Banach lattices with order continuous norms.
@@
Let $X$ be a Banach lattice with order continuous norm and let $Z$ be an ordered Banach space. Assume that there exists a positive bijection $S\in\mathcal{L}(Z,X)$, a positive operator $J\in\mathcal{L}(X,Z)$ and a sequence of positive operators $(R_{n})\subseteq\mathcal{L}(Z,X)$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to $\operatorname{id}_{Z}$ with respect to the strong operator topology.
[…]
    The space $\operatorname{span}(Z_{+})$ is a Banach lattice with order continuous norm with respect to a norm that is equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(Z_{+})}$.
[…]
    If $X$ is a KB-space, then so is $\operatorname{span}(Z_{+})$.
[…]
    If the sequence $(R_{n}J)\subseteq\mathcal{L}(X)$ converges to $\operatorname{id}_{X}$ with respect to the weak operator topology, then $J$ is a lattice homomorphism from $X$ to $\operatorname{span}(Z_{+})$ and $JX$ is a lattice ideal in $\operatorname{span}(Z_{+})$.
@@
We point out that Theorem 2.8 gives that $\operatorname{span}(Z_{+})$ is a Banach lattice if $X$ is a Banach lattice with order continuous norm, while Corollary 2.6 requires $X$ to be reflexive to get the same conclusion. Similarly, Theorem 2.8 gives that $\operatorname{span}(Z_{+})$ is a KB-space if $X$ is so, while Corollary 2.6 again requires $X$ to be reflexive to get the same conclusion. The prize that one pays for this weaker assumption in Theorem 2.8 is that one also needs the existence of the bijective positive operator $S$ and that the sequence $(JR_{n})$ is required to converge with respect to the strong instead of the weak operator topology.
@@
Indeed, after replacing sequences with subsequences, we may assume that one has $\sum_{n=1}^{\infty}\left\lVert g_{n}-g\right\rVert_{Z}<\infty$. Then the values $\left\lVert Sg_{n}-Sg\right\rVert_{X}$ are also summable, so the series $h:=\sum_{n=1}^{\infty}\left\lvert Sg_{n}-Sg\right\rvert$ converges absolutely in $X$. One gets that $0\leq Sf_{n}\leq Sg_{n}\leq h+Sg$ for all indices $n$, where the first two inequalities follow from the positivity of $S$ and the last inequality from the definition of $h$. So the sequence $(Sf_{n})$ is contained in an order interval of $X$ and since $X$ has order continuous norm, all order intervals in $X$ are weakly compact [19, Theorem 1.12]. Thus $(Sf_{n})$ has a subsequence that converges weakly in $X$. Since the inverse operator $S^{-1}:X\to Z$ (which might not be positive) is continuous by the bounded inverse theorem and is thus weakly continuous from $X$ to $Z$, it follows that $(f_{n})$ does indeed have a subsequence that converges weakly in $Z$.
Step 2: The space $\operatorname{span}(Z_{+})$ is a lattice. To see this, let $z\in\operatorname{span}(Z_{+})$. It suffices to show that ±$\pm z$ has a supremum in $\operatorname{span}(Z_{+})$. As $z\in Z_{+}-Z_{+}$, there exists $y\in Z_{+}$ such that ±$\pm z\leq y$. Then we have ±$\pm R_{n}z\leq R_{n}y$ and hence $0\leq J\left\lvert R_{n}z\right\rvert\leq JR_{n}y$ for all indices $n$. By assumption, $(JR_{n}y)$ converges to $y$ with respect to the norm in $Z$, and thus Step 1 shows that $(J\left\lvert R_{n}z\right\rvert)$ has a subsequence $(J\left\lvert R_{n_{j}}z\right\rvert)$ that converges weakly to an element $s\in Z$. As $Z_{+}$ is closed and convex it is weakly closed, so one has $s\in Z_{+}$.
We show that $s$ is the supremum of ±$\pm z$ within $\operatorname{span}(Z_{+})$. On the one hand, it follows from ±$\pm JR_{n_{j}}z\leq J\left\lvert R_{n_{j}}z\right\rvert$ for all $j$, from the convergence of $(JR_{n})$ to $\operatorname{id}_{Z}$ in the strong operator topology, and from the weak closedness of $Z_{+}$, that ±$\pm z\leq s$. On the other hand, for every upper bound $u\in\operatorname{span}(Z_{+})$ of ±$\pm z$ – repeating the above argument with $u$ instead of $y$ – we get
@@
Step 3: The space $\operatorname{span}(Z_{+})$ is a Banach lattice with order continuous norm. We first show that the cone $Z_{+}$ in $Z$ is normal. To see this, let $y,z\in Z$ satisfy $0\leq y\leq z$. Then we have $0\leq Sy\leq Sz$ in $X$ and hence, as $X$ is a Banach lattice, $\left\lVert Sy\right\rVert\leq\left\lVert Sz\right\rVert$. The inverse operator $S^{-1}$ is bounded by the bounded inverse theorem and thus,
@@
Therefore, $Z_{+}$ is normal in $Z$, and in turn, also in $\operatorname{span}(Z_{+})$; see the discussion after (1.1). So $\operatorname{span}(Z_{+})$ is an ordered Banach space with a normal cone and, at the same time, a vector lattice. According to Lemma 2.4 it is thus a Banach lattice with respect to an equivalent norm.
In order to show that the norm in $\operatorname{span}(Z_{+})$ is order continuous, let $(z_{j})$ be an increasing order bounded net in $Z_{+}$; we need to show that it is norm convergent in $\operatorname{span}(Z_{+})$. The net $(Sz_{j})$ in $X_{+}$ is also increasing and order bounded, so it is norm convergent in $X$ since $X$ has order continuous norm. By the bounded inverse theorem, we obtain that the net $(z_{j})$ is norm convergent with respect to the norm in $Z$. As $(z_{j})$ was increasing, we may apply Lemma 2.2 to assert the claim.
(b) This is essentially the same argument as in the previous paragraph – one just argues with norm-bounded instead of order-bounded nets now.
(c) As the Banach lattice $X$ has order continuous norm, it is an ideal in its bidual $X^{\prime\prime}$ [17, Theorem II.5.10] and thus, $X_{+}$ is a face in $X^{\prime\prime}_{+}$. Moreover, it follows from the convergence $JR_{n}\to\operatorname{id}_{Z}$ in the strong operator topology that $JX$ is dense in $Z$. Due to those properties and due to the convergence $R_{n}J\to\operatorname{id}_{X}$ in the weak operator topology, [3, Theorem 2.5] is applicable and gives that $JX_{+}$ is a face in $Z_{+}$ and thus in the cone of $\operatorname{span}(Z_{+})$ since the latter coincides with $Z_{+}$.
The convergence $R_{n}J\to\operatorname{id}_{X}$ in the weak operator topology also implies that $J$ is bipositive. Hence, Lemma 2.3 shows that $J$ is a lattice homomorphism from $X$ to $\operatorname{span}(Z_{+})$ and that $JX$ is a lattice ideal in $\operatorname{span}(Z_{+})$. ∎